OpenStudy (anonymous):

(ANSWER PLEASE) Let tan x = .04 and tan y = .05. Which is cot x + cot y? A. 40 B. 45 C. 4.5 D. 4.0

OpenStudy (anonymous):

\[\cot =\frac{1}{\tan}\] so \[\frac{1}{.04}+\frac{1}{.05}=? \] Simplify